Project

General

Profile

NDOS Commissioning Meeting 2011/07/07

Attendance: Jon, Susan, Jan, Nick, Dominick, Gary, Leon, Mark Mathis, Chad, Martin, Mat, Minerba

Agenda:

  1. NDOS Update
  2. DataCheck Update
  3. DAQ Update
  4. DCS Update
  5. AOB

NDOS Update

  • So far we have lost 8 of 22 golden APDs that were cooled beginning last week.
  • Tomorrow the plan is to include the rest of the golden APDs from dcm 3-1-2
  • Gary: have we learned anything from the 8 that failed? Mat: moisture was seen inside, but there was no obvious failure. Installation looked fine.
  • Peter: problem with OnlineMonitor seems to have settled down to previous failure rate ( once per ~30 minutes). Hoping someone will address this issue soon (Mark Messier perhaps)

DAQ Update

  • Group is focused on trying to get system to not require cooling to be (re)enabled after the ConfigureHardware stage.
  • Peter: status of TDUs? Andrew: TDUs installed in both MINOS and teststand, but are showing differences in timestamps. Debugging requires beam (firmware decodes specific AD signals). Could be a bad piece of hardware.
  • Peter: automatic deletion of log files is now in effect. Disk usage has gone down. 60 days worth of log files are kept on the local disk.

DataCheck Update

DCS Plans

  • Martin: yesterday, TECC temp. monitor was updated. The display looks very nice!
  • Jon: when an IOC alarms,
  • Chad: made some minor usability enhancements. Can see all dew points at once. Removed empty humidity entries. Can also print out all hard limits. Also added a "shifter tools" tab. Shifters should get kerberos ticket once/day at least.

AOB